# Analysis and Validation of Image Search Engines in Histopathology

### v1

@@ 3.3 Training/Test Data Comparison
Yottixel none TCGA (frozen) 3 pathologists
TCGA (FFPE), UPMC
@@
3.4 Comparative Experiments for Internal Data
@@
The rating $R_{i}$ for any method will simply be its position in a sorted list (proper order) based on its performance. We consistently use the F1-score of top-$n$, or the majority of top-$n$, for accuracy calculation based on leave-one-WSI-out validation.
3.4.1 Search Accuracy Results
Search accuracy is measured via $F_{1}$ Score (the harmonic mean of precision and recall). We measure the $F_{1}$ score for the top-1 search results, the majority vote among the top 3 search results (MV3), and the majority vote among the top 5 search results (MV5). Table 3 presents a comprehensive comparison of the $F_{1}$ for search accuracy for each search engine across four internal datasets Liver, Skin, CRC (Colorectal Cancer), and Breast. The Breast dataset was excluded from the MV3 and MV5 search since there are classes in this dataset with as low as one sample. We also test other variants of SISH and RetCCL (with and without post-search ranking), and Yottixel (with DenseNet, and with KimiaNet with/without post-search ranking) to provide a more comprehensive comparison.
Table 3: Average $F_{1}$ scores of top-1 accuracy (upper section), majority vote among top-3, MV@3, (middle section), and majority vote among top-5, MV@5, (lower section) on internal Mayo Clinic’s datasets. The best results are highlighted in green, the second best in blue colors.
@@ 3.4.1 Search Accuracy Results
Key Findings of Search Results Accuracy Measurements (Table 3)
@@
3.4.2 Indexing & Search Speed Experiments
Beyond accuracy, the computational expense to index and search WSIs is a key factor in selecting a capable search technology. Table 4 reports the time efficiency of the four selected search engines. It highlights two crucial aspects: the time taken to create the index for different anatomical sites ($T_{idx}$ in minutes), average indexing time per WSI ($\overline{t}_{idx}$ in minutes) and the average search times ($\overline{t}_{search}$ in minutes) for indexing and conducting the search for a query. The robustness of the search approach (in terms of number of failures) is an additional factor that we measure: being fast is not enough. The Table 4 also reports how many times each search engine failed to process a WSI.
Table 4: Index creating times $T_{idx}$ (in minutes), average indexing time per WSI $\overline{t}_{idx}$ (in minutes), average search times $\overline{t}_{search}$ (in minutes), and number of fails ($N_{\textrm{F}}$) for the internal datasets ($R_{idx}$ ranking based on indexing time, $R_{search}$) ranking based on search time, and $R_{F}$ ranking based on failures). The best results are highlighted in green, the second best in blue colors.
@@ 3.4.2 Indexing & Search Speed Experiments
Key Findings of Indexing and Search Speed Measurements (Table 4)
@@
3.4.3 Storage Requirements
Digital pathology will most likely eliminate the need for storage of glass slides but instead high-performance digital storage will be required to archive WSIs [35]. A large portion of the costs for going digital has to be spent on purchasing high-performance storage [36]. This will be a major financial factor for any hospital. The selection of the best search engine without knowing its storage requirement to index WSI archives is practically impossible. Table 5 reports the storage requirements in kilobytes (KB) for indexing and saving WSIs for different datasets. Assuming a linear relationship for all methods, the table also includes an estimate of the storage required for one million WSIs.
Table 5: Storage required to index and save each WSI [in kilobytes] for the internal datasets ($\bar{S}$ is the average index size for each search method, $\bar{S}_{10^{6}}$ is the estimate index size for one million WSIs). The best results (lowest storage requirements) are highlighted in green, the second best in blue colors.
@@ 3.4.3 Storage Requirements
Key findings of Storage Requirement (Table 5)
@@
Perhaps the most important measure for the storage (and naturally for the speed) of each search method is the average index size per WSI. With 0.03 kilobytes per WSI, BoVW is the most efficient search approach. SISH needs $\approx$98 kilobytes per WSI which clearly is a tremendous overhead for any digital archive.
3.5 Comparative Experiments for External Data
@@
3.5.1 Search Accuracy Results
For further evaluation of the search engines, we apply them on three well-known WSI-level public datasets: CAMELYON16 [37] (short CAM16), BRACS [38], and PANDA [39]. As mentioned before, due to the extensive nature of leave-one-out validation, we restricted these experiments only using MV@5 and only to SISH and Yottixel as BoVW and RetCCL displayed low accuracy. Table 6 shows the results. Yottixel-K is the most accurate method for CAM16 and BRACS. It is also the best method with an average $F_{1}$ score of 56.3%. SISH lands in the second place with an average $F_{1}$ score of 52.9%. However, as Table 7 shows, SISH fails in 57% of cases (for 569/1000 WSIs) in the PANDA dataset.
Table 6: Results for MV@5 on the public datasets: CAM16, BRACS, and PANDA.
@@ 3.5.1 Search Accuracy Results
3.5.2 Search Speed Benchmark
Table 7 reports the time efficiency of the selected search engines. It highlights two crucial aspects: the time taken to create the index for different anatomical sites ($T_{idx}$ in minutes) and the average search times ($T_{search}$ in seconds) for conducting searches. It also demonstrates how many times each search engine failed to process a WSI among datasets. Note, the search time encompasses both the indexing time and the matching time. Table 7 confirms what we reported for our internal (private) dataset in Table 4. SISH is much slower and fails more often than Yottixel.
Table 7: Index creating times $T_{idx}$ (in minutes), average indexing time per WSI $\overline{t}_{idx}$(in minutes), average search times $\overline{t}_{search}$ (in minutes), and number of fails ($N_{\textrm{F}}$) for public dataset.
@@ 3.5.2 Search Speed Benchmark
3.5.3 Storage Benchmark
Table 8 presents storage requirements in kilobytes (KB) for indexing and saving WSIs for different datasets. It lists the storage requirements for various methods across all datasets. The table also includes an estimate of the storage required for one million WSIs. The results for the three public datasets are aligned with the results we reported for our internal (private) datasets in Table 8. The indexing overhead of SISH is prohibitively large (165 kilobytes per WSI) whereas Yottixel is rather lean (0.23 kilobytes per WSI).
Table 8: Storage required to index and save each WSI [in kilobytes] for the external datasets ($\bar{S}$ is the average index size per dataset, $\bar{S}_{10^{6}}$ is the estimate index size for one million WSIs). The best results (lowest storage requirements) are highlighted in green, the second best in blue colors.
@@ 3.5.3 Storage Benchmark
3.6 Overall Performance Ranking for Internal & External Data
@@
[…]
@@
Table 9: Overall Performance for Internal Datasets: Based on all previous ratings (for top-1, MV3, MV5, indexing time, search time, number of failed WSIs, and required storage), we calculate $R^{*}_{\textrm{internal}}$ as the average performance rank for each search method.
@@
Yottixel 3 2 1 2 1 1 2 1.71
Yottixel-KR 2 1 2 2 1 2 2 1.71
Yottixel-K 1 3 4 2 1 2 2 2.14
@@
SISH 4 4 3 3 3 4 4 3.57
SISH-N 5 5 5 3 3 4 4 4.14
RetCCL-N 8 8 7 4 4 3 3 5.28
RetCCL 7 7 8 4 4 3 3 5.43
@@
[…]
Table 10 shows the same ratings for the external dataset. As the two more accurate methods (compared to BoVW and RetCCL), Yottixel and SISH were only tested for MV5. The results confirm the general rating that Yottixel is more accurate, more efficient, and more reliable than SISH. The only difference with internal results is that Yottixel-K climbs to the first place ahead of Yottixel. This means that even for a method that is generally reliable, the selection of a deep network may still need some trial-and-error for specific datasets.
[…]
@@ 4 Discussions
As previously discussed, the critical task of patient matching through matching WSI to other WSIs underscores the need for an efficient generation of WSI representations (like Yottixel’s mosaic) and a reliable scheme for WSI-to-WSI comparison (like Yottixel’s median-of-minimum distance calculation). This is where SISH and RetCCL fall short, as they not only do not possess the capability to generate expressive WSI representations, but also they lose the capability of WSI-to-WSI matching. The former shortcoming for SISH is due to crippling Yottixel’s barcodes with a patch-oriented tree structure and a post-search ranking scheme and for RetCCL due to the adoption of the SISH ranking scheme. The latter shortcoming is mainly due to using a redundant ranking algorithm after the search.
@@
$\blacksquare$ RetCCL barely qualifies as a search engine. Its primary focus is on the CCL network, which appears incapable of providing expressive embeddings for tissue morphology.
@@
$\blacksquare$ The accuracy level of all search engines is still low for clinical utility.
$\blacksquare$ We should prioritize both speed and storage considerations in digital pathology.
$\blacksquare$ No automated curation framework exists for WSI selection.
$\blacksquare$ No algorithm has been put forward for setting magnification level and patch size.
$\blacksquare$ No multimodal search approach has been proposed.
5 Methods
This section describes the process of our experiments, how the search engines were implemented, what pre-trained networks were involved, and what dataset was used.
5.1 Experimental Setup
Search accuracy was measured among the top-1, majority vote among top-3, and majority vote among top-5 WSI matching results. The $F_{1}$ score was used as a measure to quantify both precision and recall. Additionally, the macro-averaged accuracy ($\bar{F}_{1}$) was calculated for each search engine in each experiment. All the experiments regarding SISH, Yottixel, and RetCCL were carried out using a leave-one-patient-out approach which is the extreme form of cross-validation where the number of folds is equal to the number of patients in the dataset. Since the BoVW approach requires training an autoencoder for each dataset, the experiments for this specific search engine were carried out via a 5-fold cross-validation. We also measured the speed of each search engine in indexing and searching, separately. The average time taken to create the index was reported in minutes and the average search times (including both indexing and matching time for the query) were reported in seconds. To quantify reliability, we also counted how many times each search engine failed to process a WSI in each dataset. The failed WSIs were excluded from that specific experiment of search engine and dataset. We also measured the storage need of each search engine, separately. After indexing all the WSIs in each dataset, we measured the total disk space each search engine occupied to index the WSIs, excluding failed ones, to obtain the average disk space needed for each WSI. We also extrapolated the estimated amount of disk space needed to index one million WSIs using the estimated average.
5.2 Implementation
Yottixel. We used the sample Python code publicly published on KimaLab’s website [^2] https://kimialab.uwaterloo.ca/kimia/index.php/yottixel/. The published code demonstrates the steps needed to index and search one sample WSI. We used the exact code to index all the WSIs iteratively. We then utilized the saved indexed images to search for a query WSI. The sample code uses a pre-trained DenseNet121 for feature extraction which is the original proposed version of Yottixel. As previously explained in the experiment setup section, we also tested two new variants of the original Yottixel. One variant was created by swapping the DenseNet121 with KimiaNet as the feature extractor. The other variant was implemented to use post-search ranking, as it is already implemented in SISH. To do so, we got the patch-level features extracted by Yottixel and applied the same ranking used by SISH. It is noteworthy that this change rendered Yottixel unable to match WSI-2-WSI, as this is already the case for SISH.
5.3 Pre-Trained Networks
Besides the deep networks published as part of the search engine methods, we used two other pre-trained networks as the backbone, as needed. One is DenseNet121 [41] which is a general-purpose CNN and the other is KimiaNet, a pathology-specific DenseNet.
KimiaNet is a deep network fine-tuned for histopathology image representation. It is a CNN based on the DenseNet-121 architecture with four dense blocks, fine-tuned and trained with histopathology images from TCGA WSIs. It is trained using more than 240,000 image patches acquired at 20X magnification from 7126 whole slide images of formalin-fixed paraffin-embedded human pathology samples [40].
5.4 Datasets
We conducted our benchmarking on four internal and three public datasets. The internal datasets are collected from the Mayo Clinic each from a different organ with different labels: breast epithelial tumor subtypes, fatty liver disease, cutaneous squamous cell carcinoma (cSCC) subtypes, and colorectal polyps. Table 11 gives a detailed overview of the number of patients, WSIs, and classes in these public/internal datasets.
Table 11: Validation Datasets. The number of patients, WSIs, and classes in each dataset used for benchmarking the search engines.
Internal Dataset Number of Patients Number of WSIs Number of Classes
Breast Epithelial Tumors 74 74 16 + normal
Fatty Liver Disease 326 326 2 + normal
cSCC 660 660 3 + normal
Colorectal Polyps 209 209 3
Public CAM16 [37] 129 129 2
BRACS [38] 87 87 3
PANDA [39] 1000 1000 6
  • cSCC = cutaneous squamous cell carcinoma
5.4.1 Internal Datasets
5.4.2 Public Datasets
@@
Part of the code used in the current study has already been published publicly in various locations. These code segments are openly accessible through their respective sources. The complete set of codes employed for benchmarking purposes will be made publicly available upon the publication of the paper.

### v2

@@ IV-C Analysis of Training/Test Data
Yottixel none TCGA (frozen) 3
TCGA (FFPE)
UPMC
@@
IV-D Comparative Experiments for Internal Data
@@
The rating $R_{i}$ for any method will simply be its position in a sorted list (proper order) based on its performance. We consistently use the F1-score of top-$n$, or the majority of top-$n$, for accuracy calculation based on leave-one-WSI-out validation. The F1-score is the harmonic mean of precision (number of the retrieved images actually relevant to the query (i.e., low false-positive rate) and recall (number of images with the same class label as the query were actually retrieved, i.e., low false-negative rate), providing a balanced measure of both metrics [59].
IV-D1 Search Accuracy Results
Search accuracy is measured via $F_{1}$ Score (the harmonic mean of precision and recall). We measure the $F_{1}$ score for the top-1 search results, the majority vote among the top 3 search results (MV3), and the majority vote among the top 5 search results (MV5). Table V presents a comprehensive comparison of the $F_{1}$ for search accuracy for each search engine across four internal datasets Liver, Skin, CRC (Colorectal Cancer), and Breast. The Breast dataset was excluded from the MV3 and MV5 search since there are classes in this dataset with as low as one sample. We also test other variants of SISH and RetCCL (with and without post-search ranking), and Yottixel (with DenseNet, and with KimiaNet with/without post-search ranking) to provide a more comprehensive comparison.
TABLE V: Average $F_{1}$ scores of top-1 accuracy (upper section), majority vote among top-3, MV@3, (middle section), and majority vote among top-5, MV@5, (lower section) on internal Mayo Clinic’s datasets. The best results are highlighted in green, the second best in blue colors.
@@ IV-D1 Search Accuracy Results
Key Findings of Search Results Accuracy Measurements (Table V)
@@
IV-D2 Indexing & Search Speed Experiments
Beyond accuracy, the computational expense to index and search WSIs is a key factor in selecting a capable search technology. Table VI reports the time efficiency of the four selected search engines. It highlights two crucial aspects: the time taken to create the index for different anatomical sites ($T_{idx}$ in minutes), average indexing time per WSI ($\overline{t}_{idx}$ in minutes) and the average search times ($\overline{t}_{search}$ in minutes) for indexing and conducting the search for a query. The robustness of the search approach (in terms of number of failures) is an additional factor that we measure: being fast is not enough. The Table VI also reports how many times each search engine failed to process a WSI. The predominant reason for failures appears to be absence of any patch for processing after pre- or post-processing of patches.
TABLE VI: Index creating times $T_{idx}$ (in minutes), average indexing time per WSI $\overline{t}_{idx}$ (in minutes), average search times $\overline{t}_{search}$ (in minutes), and number of fails ($N_{\textrm{F}}$) for the internal datasets ($R_{idx}$ ranking based on indexing time, $R_{search}$) ranking based on search time, and $R_{F}$ ranking based on failures). The best results are highlighted in green, the second best in blue colors.
@@ IV-D2 Indexing & Search Speed Experiments
Key Findings of Indexing and Search Speed Measurements (Table VI)
@@
IV-D3 Storage Requirements
Digital pathology no longer relies on the examination of glass slides; however, it necessitates high-performance digital storage to effectively archive gigapixel WSIs[65]. A significant portion of the expenses associated with transitioning to a digital system is allocated to the acquisition of high-performance storage solutions [66]. This will be a major financial factor for any hospital. The selection of the best search engine without knowing its storage requirement to index WSI archives is practically impossible. Table VII reports the storage requirements in kilobytes (KB) for indexing and saving WSIs for different datasets. Assuming a linear relationship for all methods, the table also includes an estimate of the storage required for one million WSIs. It’s important to consider that the stored index needs to be loaded into memory for the search & matching to take place.
TABLE VII: Storage required to index and save each WSI [in kilobytes] for the internal datasets ($\bar{S}$ is the average index size for each search method, $\bar{S}_{10^{6}}$ is the estimate index size for one million WSIs). The best results (lowest storage requirements) are highlighted in green, the second best in blue colors.
@@ IV-D3 Storage Requirements
Key findings of Storage Requirement (Table VII)
@@
Perhaps the most important measure for the storage (and naturally for the speed) of each search method is the average index size per WSI. With 0.03 kilobytes per WSI, BoVW is the most efficient search approach. SISH needs $\approx$98 kilobytes per WSI which clearly is a tremendous overhead for any digital archive, and impractical for loading into memory to conduct search.
IV-E Comparative Experiments for External Data
@@
IV-E1 Search Accuracy Results
For further evaluation of the search engines, we apply them on three well-known WSI-level public datasets: CAMELYON16 [56] (short CAM16), BRACS [57], and PANDA [58]. As mentioned before, due to the extensive nature of leave-one-out validation, we restricted these experiments only using MV@5 and only to SISH and Yottixel as BoVW and RetCCL displayed low accuracy. Table VIII shows the results. Yottixel-K is the most accurate method for CAM16 and BRACS. It is also the best method with an average $F_{1}$ score of 56.3%. SISH lands in the second place with an average $F_{1}$ score of 52.9%. However, as Table IX shows, SISH fails in 57% of cases (for 569/1000 WSIs) in the PANDA dataset.
TABLE VIII: Results for MV@5 on the public datasets: CAM16, BRACS, and PANDA.
@@ IV-E1 Search Accuracy Results
Yottixel 64.5$\pm$19.1 52.6$\pm$9.0 31.8$\pm$12.4 49.6$\pm$16.5 3
SISH 73.5$\pm$13.4 49.0$\pm$8.9 36.2$\pm$20.2 $\ddagger$ 52.9$\pm$19.0 $\ddagger$ 2
Yottixel-K 75.5$\pm$13.4 59.7$\pm$4.7 33.8$\pm$12.0 56.3$\pm$21.0 1
Yottixel-KR 73.5$\pm$13.4 51.7$\pm$5.0 35.5$\pm$11.0 45.0$\pm$8.50 5
SISH-N 59.0$\pm$28.3 51.0$\pm$5.0 29.0$\pm$14.6 46.3$\pm$15.5 4
  • $\ddagger$
    SISH failed in search for almost 57% of PANDA images (see Table IX)
IV-E2 Search Speed Benchmark
Table IX reports the time efficiency of the selected search engines. It highlights two crucial aspects: the time taken to create the index for different anatomical sites ($T_{idx}$ in minutes) and the average search times ($T_{search}$ in seconds) for conducting searches. It also demonstrates how many times each search engine failed to process a WSI among datasets. Note, the search time encompasses both the indexing time and the matching time. Table IX confirms what we reported for our internal (private) dataset in Table VI. SISH is much slower and fails more often than Yottixel.
TABLE IX: Index creating times $T_{idx}$ (in minutes), average indexing time per WSI $\overline{t}_{idx}$(in minutes), average search times $\overline{t}_{search}$ (in minutes), and number of fails ($N_{\textrm{F}}$) for public dataset.
@@ IV-E2 Search Speed Benchmark
IV-E3 Storage Benchmark
Table X presents storage requirements in kilobytes (KB) for indexing and saving WSIs for different datasets. It lists the storage requirements for various methods across all datasets. The table also includes an estimate of the storage required for one million WSIs. The results for the three public datasets are aligned with the results we reported for our internal (private) datasets in Table X. The indexing overhead of SISH is prohibitively large (165 kilobytes per WSI) whereas Yottixel is rather lean (0.23 kilobytes per WSI).
TABLE X: Storage required to index and save each WSI [in kilobytes] for the external datasets ($\bar{S}$ is the average index size per dataset, $\bar{S}_{10^{6}}$ is the estimate index size for one million WSIs). The best results (lowest storage requirements) are highlighted in green, the second best in blue colors.
@@ IV-E3 Storage Benchmark
IV-F Overall Performance Ranking for Internal & External Data
@@
Table XI presents an average performance rank for internal datasets, denoted as $R$ based on these individual SER values. Each row corresponds to a specific search engine or method, such as Yottixel-KR, Yottixel, Yottixel-K, BoVW, SISH-N, SISH, RetCCL-N, and RetCCL. The table provides rankings for each search method across the various R metrics. The $R$ values (see equation 1) provide a composite ranking, allowing readers to identify which method performs better overall, taking into account a range of performance measures.
@@
TABLE XI: Overall Performance for Internal Datasets: Based on all previous ratings (for top-1, MV3, MV5, indexing time, search time, number of failed WSIs, and required storage), we calculate $R^{*}_{\textrm{internal}}$ as the average performance rank for each search method.
@@
Yottixel $\ddagger$ 3 2 1 2 1 1 2 1.71
Yottixel-KR 2 2 1 2 2 1 2 2 1.71
Yottixel-K 1 1 3 4 2 1 2 2 2.14
@@
SISH $\ddagger$ 4 4 3 3 3 4 4 3.57
SISH-N 4 5 5 5 3 3 4 4 4.14
RetCCL-N 5 8 8 7 4 4 3 3 5.28
RetCCL $\ddagger$ 7 7 8 4 4 3 3 5.43
@@
  • SER = search engine ranking in each benchmark
Table XII shows the same ratings for the external dataset. As the two more accurate methods (compared to BoVW and RetCCL), Yottixel and SISH were only tested for MV5. The results confirm the general rating that Yottixel is more accurate, more efficient, and more reliable than SISH. The only difference with internal results is that Yottixel-K climbs to the first place ahead of Yottixel. This means that even for a method that is generally reliable, the selection of a deep network may still need some trial-and-error for specific datasets.
TABLE XII: Overall Performance for External Datasets: Based on all previous SER (for top1, MV@3, MV@5, indexing time, search time, number of failed WSIs, and required storage), we calculate $R$ as the average performance rank for each search method.
@@ V Discussions
The impact of deep features on search accuracy is expected to be high. We know that stain variations [67, 68, 69] and data bias [70, 71] may affect the equality of embeddings. One may design and train a general model feature extraction in histopathology. For instance, Virchow model used one million WSIs for its training [72]. As well, one might think in direction of customized training customized networks for specific organs and diseases [73, 59].
@@
As previously discussed, the critical task of patient tissue matching through matching WSI to other WSIs underscores the need for an efficient generation of WSI representations (like Yottixel’s mosaic) and a reliable scheme for WSI-to-WSI comparison (like Yottixel’s median-of-minimum distance calculation). This is where SISH and RetCCL fall short, as they not only do not possess the capability to generate expressive WSI representations, but also they lose the capability of WSI-to-WSI matching. The former shortcoming for SISH is due to crippling Yottixel’s barcodes with a patch-oriented tree structure and a post-search ranking scheme and for RetCCL due to the adoption of the SISH ranking scheme. The latter shortcoming is mainly due to using a redundant ranking algorithm after the search. Both the vEB tree and the post-search ranking separate mosaic patches, rendering mosaic-to-mosaic comparison infeasible.
@@
$\blacksquare$ RetCCL barely qualifies as a search engine. Its primary focus is on the CCL network, which appears to provide non-expressive embeddings for tissue morphology when used for image retrieval.
$\blacksquare$ All four methods can incrementally index new WSIs although they may then perform poor if the new tissue pattern does not fall inside the latent space of the deep network used.
@@
  • The accuracy level of all search engines is still low for clinical utility.
  • We should prioritize both speed and storage considerations in digital pathology.
  • No automated curation framework exists for WSI selection.
  • No algorithm has been put forward for setting magnification level and patch size.
  • No multimodal search approach has been proposed.
@@
The Yottixel’s code, and KimiaNet model and its weights are available at Kimia Lab’s GitHub (https://github.com/KimiaLabMayo). SISH code is available at a GitHub repository (https://github.com/mahmoodlab/SISH). RetCCL code and model are available at a GitHub repository (https://github.com/Xiyue-Wang/RetCCL). BoVW code is available at Kimia Lab’s GitHub (https://github.com/KimiaLabMayo).
Disclosures
Hamid R. Tizhoosh is the inventor of Yottixel’s barcoding, a patented technology that is owned by Huron Digital Pathology (St. Jacobs, ON, Canada) based on sponsorship agreements. Hamid R. Tizhoosh has no relationship with Huron Digital Pathology since December 2020.